Spud, Inc. a manufacturer of gourmet potato chips, employs activity-based costing. The budgeted data for each of the activity cost pools is provided below for the year 2020. Activity Cost Pools Estimated Overhead Estimated Use of Cost Drivers per Activity Ordering and receiving $93,572 14,900 orders Food processing 522,660 62,000 machine hours Packaging 1,927,200 440,000 labor hours For 2020, the company had 10,200 orders and used 51,000 machine hours, and labor hours totaled 509,000. Correct answer. Your answer is correct. Calculate the overhead rates for each activity. (Round answers to 2 decimal places, e.g. 12.25.) Overhead Rates Ordering and receiving $Entry field with correct answer per order Food processing $Entry field with correct answer per machine hour Packaging $Entry field with correct answer per labor hour SHOW SOLUTION SHOW ANSWER LINK TO TEXT Incorrect answer. Your answer is incorrect. Try again. What is the total overhead applied? Total overhead applied $Entry field with incorrect answer
Answer #1
1
Overhead Rates
Ordering and receiving 6.28 per order =93572/14900
Food processing 8.43 per machine hour =522660/62000
Packaging 4.38 per labor hour =1927200/440000
2
Ordering and receiving 64056 =10200*6.28
Food processing 429930 =51000*8.43
Packaging 2229420 =509000*4.38
Total overhead applied 2723406